# TimeDiT: General-purpose Diffusion Transformers for Time Series Foundation Model

### v1

@@ 6 Conclusion
In this paper, we introduced TimeDiT, a pioneering approach to creating a versatile and robust foundation model for various time series tasks under practical scenarios. By integrating transformer architecture with diffusion model, TimeDiT effectively captures temporal dependencies and addresses real world challenges unique to time series regarding multi-resolution and missing values as well as incorporating external knowledge. Our innovative masking strategies allow for a consistent training framework adaptable to diverse tasks such as forecasting, imputation, and anomaly detection and synthetic data generation. Extensive experiments demonstrated the strong performance of TimeDiT on both practical scenarios and standard benchmarks. However, we recognize some limitations. We primarily explored common sequence lengths and did not assess TimeDiT’s performance on very long sequences. While we have introduced randomness in prediction length and feature numbers up to a maximum, we aim to develop more scalable solutions for highly variable multivariate time series. Additionally, our understanding of how different types of external information contribute to performance is still developing. For future work, we envision several key directions: enhancing scalability to improve TimeDiT’s ability to handle practical time series with varying multivariate numbers; developing techniques for seamless multi-modal integration, allowing TimeDiT to leverage diverse data sources for improved performance across different tasks; and extending TimeDiT’s capabilities to effectively process and analyze very long time series sequences, addressing a critical need in many real-world applications.
Appendix A Datasets
@@
    The ETT datasets Zhou et al. (2021) [^1] ETT: https://github.com/zhouhaoyi/ETDataset include electricity load data at various resolutions (ETTh & ETTm) from two different electricity stations.
@@
    The Weather dataset comprises 21 meteorological indicators collected in Germany over the span of one year.
@@
    The Electricity dataset provides information on electricity consumption.
@@
    The SMD dataset Su et al. (2019) includes multivariate time-series data collected from server machines in a data center. It typically contains metrics such as CPU usage, memory usage, and disk activity.
@@
    The PSM dataset Abdulaal et al. (2021)is used for predictive maintenance and includes sensor data from industrial machines. It often contains readings such as temperature, pressure, and vibration over time.
@@
    The MSL dataset Hundman et al. (2018) comes from the Mars Science Laboratory mission, specifically the Curiosity rover. It includes telemetry data from the rover’s sensors and systems.
@@
    The SWaT dataset Mathur and Tippenhauer (2016) originates from a scaled-down water treatment testbed designed to reflect a real-world water treatment process. It includes sensor and actuator data collected over time.
@@
    The SMAP dataset Hundman et al. (2018)comes from NASA’s Soil Moisture Active Passive (SMAP) mission, which measures soil moisture and freeze/thaw state. It includes time-series data from multiple sensors aboard the SMAP satellite.
@@
    The Sine dataset Yoon et al. (2019) is synthetically generated by sinusoidal waves.
@@
    The Air Quality dataset  [^2] Air Quality: https://archive.ics.uci.edu/dataset/360/air+qualitycontains hourly averaged readings from five metal oxide chemical sensors integrated into an Air Quality Chemical Multisensor Device. This device was positioned at road level in a highly polluted area of an Italian city. Data were collected from March 2004 to February 2005, making it the longest freely available record of on-field air quality chemical sensor responses.
@@
    The Stock dataset  [^3] Stock: https://finance.yahoo.com/quote/GOOG contains daily historical Google stocks data from 2004 to 2019.
@@
    The UCI Appliances Energy prediction dataset  [^4] Energy: https://archive.ics.uci.edu/ml/datasetsconsists of multivariate, continuous-valued measurements including numerous temporal features measured at close intervals.
@@
    The Cloud dataset: The Huawei cloud datasets contain serverless traces (Joosen et al., 2023). Following (Rasul et al., 2023), we selected 8 time series containing metrics based on the minute-frequency occurrences of the top 10 functions over a period of 141 days. The metrics included in these series are: Function delay; Platform delay; CPU usage; Memory usage; CPU limit; Memory limit; Instances; Requests. The functions were chosen based on their median occurrences throughout the dataset.
@@
    The Weather_2 dataset: The Weather_2 dataset comprises hourly climate time series data collected near Monash University, Clayton, Victoria, Australia, from January 2010 to May 2021. It includes series for temperature, dewpoint temperature, wind speed, mean sea level pressure, relative humidity, surface solar radiation, surface thermal radiation, and total cloud cover (Godahewa et al., 2021).
@@
Appendix B Forecasting Experiment Setting
For the forecasting task, we utilized five widely-used open datasets to evaluate probabilistic time series forecasting performance. These datasets were collected in GluonTS  (Alexandrov et al., 2020) and have been previously employed in  (Tashiro et al., 2021; Salinas et al., 2019):
Table  8 summarizes the characteristics of each dataset. The task for these datasets is to predict the future $L_{2}$ steps given the observed $L_{1}$ steps. We set $L_{1}$ and $L_{2}$ values based on previous studies  (Tashiro et al., 2021; Salinas et al., 2019). For training, we randomly selected $L_{1}+L_{2}$ consecutive time steps as a single time series and designated the last $L_{2}$ steps as forecasting targets. We adhered to the train/test splits used in previous studies and utilized the last five samples of the training data as validation data.
For the full-shot setting, we trained separate models on different datasets. Due to the large number of features in multivariate time series, we adopted subset sampling of features for training. For each input, we split them into subsets based on their order. If the last subset was smaller than the fixed shape, we applied padding to ensure equal input sizes across all subsets. In the multi-resolution setting, we used different resolutions identified by the resolution number, which corresponded to different sampling rates for the exchange features. It is worth to note that the aforementioned strategy was also employed for zero-shot training, as the input feature length varied across datasets.
Appendix C Training details
The codebase for TimeDiT is modified from https://github.com/facebookresearch/DiT, where they provide different model sizes, including Small (S), Big (B), Large (L), Extra Large (XL) (Peebles and Xie, 2022). In our training, we used Adam optimizer with a training rate of 0.0001 without weight decay. Batch size is set to 512. The maximum channel number $K_{max}$ is set to 40. All experiments are run on NVIDIA A100 GPUs. The zero-shot foundation model was trained on the ETT, weather, illness, air quality, and cloud datasets and used for different downstream tasks. We will include more available time series datasets to develop a more robust time series foundation model as the future work.
Table 8: Evaluate time series dataset for forecasting tasks.
Solar 137 7009 168 24 7 1 hour $\mathbb{R}^{+}$
Electricity 370 5833 168 24 7 1 hour $\mathbb{R}^{+}$
Traffic 963 4001 168 24 7 1 hour (0, 1)
Taxi 1214 1488 48 24 56 30 mins $\mathbb{N}$
Exchange 8 6071 60 30 5 1 day $\mathbb{R}^{+}$
Table 9: Training Details. Imp stands for Imputation. SG stands for Syntheric Generation. AD stands for Anomaly Detection. FC stands for Forecasting
Dataset Task Loss model size hidden size attention head depth
ETTh Imp $L_{simple}$ S 384 6 12
ETTm Imp $L_{simple}$ S 384 6 12
Weather Imp $L_{simple}$ S 384 6 12
Electricity Imp $L_{simple}$ S 384 6 12
Air Quality Imp $L$ S 384 6 12
Sine SG $L$ S 384 6 12
Stock SG $L$ S 384 6 12
Energy SG $L$ S 384 6 12
MSL AD $L_{simple}$ S 384 6 12
PSM AD $L_{simple}$ S 384 6 12
SMAP AD $L_{simple}$ S 384 6 12
SMD AD $L_{simple}$ S 384 6 12
SWaT AD $L_{simple}$ S 384 6 12
Solar FC $L_{simple}$ B 768 12 12
Taxi FC $L_{simple}$ B 768 12 12
Traffic FC $L_{simple}$ B 768 12 12
Exchange FC $L_{simple}$ B 768 12 12
Electricity FC $L_{simple}$ B 768 12 12
Appendix D Metrics
@@
$\text{CRPS}(F^{-1},x)\approx\frac{1}{19}\sum_{i=1}^{19}2\Lambda_{i\cdot 0.05}(F^{-1}(i\cdot 0.05),x).$
@@
$\frac{\sum_{k,l}\text{CRPS}(F^{-1}_{k,l},x_{k,l})}{\sum_{k,l}|x_{k,l}|}$
where $k$ and $l$ denote the features and time steps of the imputation targets, respectively.
CRPSsum measures CRPS for the distribution $F$ of the sum of all $K$ features, calculated by:
$\frac{\sum_{l}\text{CRPS}(F^{-1},\sum_{k}x_{k,l})}{\sum_{k,l}|x_{k,l}|}$
@@
Appendix E Synthetic Generation
We use 80% of all data for training and evaluate on the same data. For the air quality dataset, previous methods did not carefully the -200 values as placeholder for missing values. In our experiment, we masked all the -200 values for TimeDiT and baselines that support masks. For baselines that do not support mask, we replace -200 with the mean value. Minmax scaler is used for all models. Figure 4, 3,5,6 shows the PCA plots for all datasets and baselines. The visual comparison also validates the superiority of TimeDiT.
Figure 4: PCA plot for sine dataset.
Figure 5: PCA plot for air quality dataset.
Figure 6: PCA plot for energy dataset.
Appendix F Imputation
Methods TimeDiT GPT2(3) TimesNet PatchTST ETSformer LightTS DLinear FEDformer Stationary Autoformer Informer Reformer
Mask Ratio MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$ETTh1$ 12.5% 0.025 0.101 0.043 0.140 0.057 0.159 0.093 0.201 0.126 0.263 0.240 0.345 0.151 0.267 0.070 0.190 0.060 0.165 0.074 0.182 0.114 0.234 0.074 0.194
25% 0.034 0.122 0.054 0.156 0.069 0.178 0.107 0.217 0.169 0.304 0.265 0.364 0.180 0.292 0.106 0.236 0.080 0.189 0.090 0.203 0.140 0.262 0.102 0.227
37.5% 0.047 0.143 0.072 0.180 0.084 0.196 0.120 0.230 0.220 0.347 0.296 0.382 0.215 0.318 0.124 0.258 0.102 0.212 0.109 0.222 0.174 0.293 0.135 0.261
50% 0.063 0.166 0.107 0.216 0.102 0.215 0.141 0.248 0.293 0.402 0.334 0.404 0.257 0.347 0.165 0.299 0.133 0.240 0.137 0.248 0.215 0.325 0.179 0.298
Avg 0.042 0.135 0.069 0.173 0.078 0.187 0.115 0.224 0.202 0.329 0.284 0.373 0.201 0.306 0.117 0.246 0.094 0.201 0.103 0.214 0.161 0.279 0.122 0.245
$ETTh2$ 12.5% 0.025 0.107 0.039 0.125 0.040 0.130 0.057 0.152 0.187 0.319 0.101 0.231 0.100 0.216 0.095 0.212 0.042 0.133 0.044 0.138 0.305 0.431 0.163 0.289
25% 0.037 0.129 0.044 0.135 0.046 0.141 0.061 0.158 0.279 0.390 0.115 0.246 0.127 0.247 0.137 0.258 0.049 0.147 0.050 0.149 0.322 0.444 0.206 0.331
37.5% 0.046 0.149 0.051 0.147 0.052 0.151 0.067 0.166 0.400 0.465 0.126 0.257 0.158 0.276 0.187 0.304 0.056 0.158 0.060 0.163 0.353 0.462 0.252 0.370
50% 0.062 0.173 0.059 0.158 0.060 0.162 0.073 0.174 0.602 0.572 0.136 0.268 0.183 0.299 0.232 0.341 0.065 0.170 0.068 0.173 0.369 0.472 0.316 0.419
Avg 0.042 0.139 0.048 0.141 0.049 0.146 0.065 0.163 0.367 0.436 0.119 0.250 0.142 0.259 0.163 0.279 0.053 0.152 0.055 0.156 0.337 0.452 0.234 0.352
$ETTm1$ 12.5% 0.016 0.083 0.017 0.085 0.023 0.101 0.041 0.130 0.096 0.229 0.093 0.206 0.080 0.193 0.052 0.166 0.032 0.119 0.046 0.144 0.063 0.180 0.042 0.146
25% 0.019 0.091 0.022 0.096 0.023 0.101 0.044 0.135 0.096 0.229 0.093 0.206 0.080 0.193 0.052 0.166 0.032 0.119 0.046 0.144 0.063 0.180 0.042 0.146
37.5% 0.025 0.102 0.029 0.111 0.029 0.111 0.049 0.143 0.133 0.271 0.113 0.231 0.103 0.219 0.069 0.191 0.039 0.131 0.057 0.161 0.079 0.200 0.063 0.182
50% 0.032 0.115 0.040 0.128 0.036 0.124 0.055 0.151 0.186 0.323 0.134 0.255 0.132 0.248 0.089 0.218 0.047 0.145 0.067 0.174 0.093 0.218 0.082 0.208
Avg 0.023 0.098 0.028 0.105 0.027 0.107 0.047 0.140 0.120 0.253 0.104 0.218 0.093 0.206 0.062 0.177 0.036 0.126 0.051 0.150 0.071 0.188 0.055 0.166
$ETTm2$ 12.5% 0.016 0.065 0.017 0.076 0.018 0.080 0.026 0.094 0.108 0.239 0.034 0.127 0.062 0.166 0.056 0.159 0.021 0.088 0.023 0.092 0.133 0.270 0.108 0.228
25% 0.022 0.078 0.020 0.080 0.020 0.085 0.028 0.099 0.164 0.294 0.042 0.143 0.085 0.196 0.080 0.195 0.024 0.096 0.026 0.101 0.135 0.272 0.136 0.262
37.5% 0.027 0.089 0.022 0.087 0.023 0.091 0.030 0.104 0.237 0.356 0.051 0.159 0.106 0.222 0.110 0.231 0.027 0.103 0.030 0.108 0.155 0.293 0.175 0.300
50% 0.031 0.099 0.025 0.095 0.026 0.098 0.034 0.110 0.323 0.421 0.059 0.174 0.131 0.247 0.156 0.276 0.030 0.108 0.035 0.119 0.200 0.333 0.211 0.329
Avg 0.024 0.083 0.021 0.084 0.022 0.088 0.029 0.102 0.208 0.327 0.046 0.151 0.096 0.208 0.101 0.215 0.026 0.099 0.029 0.105 0.156 0.292 0.157 0.280
$ECL$ 12.5% 0.051 0.148 0.080 0.194 0.085 0.202 0.055 0.160 0.196 0.321 0.102 0.229 0.092 0.214 0.107 0.237 0.093 0.210 0.089 0.210 0.218 0.326 0.190 0.308
25% 0.061 0.163 0.087 0.203 0.089 0.206 0.065 0.175 0.207 0.332 0.121 0.252 0.118 0.247 0.120 0.251 0.097 0.214 0.096 0.220 0.219 0.326 0.197 0.312
37.5% 0.074 0.181 0.094 0.211 0.094 0.213 0.076 0.189 0.219 0.344 0.141 0.273 0.144 0.276 0.136 0.266 0.102 0.220 0.104 0.229 0.222 0.328 0.203 0.315
50% 0.090 0.202 0.101 0.220 0.100 0.221 0.091 0.208 0.235 0.357 0.160 0.293 0.175 0.305 0.158 0.284 0.108 0.228 0.113 0.239 0.228 0.331 0.210 0.319
Avg 0.069 0.174 0.090 0.207 0.092 0.210 0.072 0.183 0.214 0.339 0.131 0.262 0.132 0.260 0.130 0.259 0.100 0.218 0.101 0.225 0.222 0.328 0.200 0.313
$Weather$ 12.5% 0.029 0.033 0.026 0.049 0.025 0.045 0.029 0.049 0.057 0.141 0.047 0.101 0.039 0.084 0.041 0.107 0.027 0.051 0.026 0.047 0.037 0.093 0.031 0.076
25% 0.031 0.033 0.028 0.052 0.029 0.052 0.031 0.053 0.065 0.155 0.052 0.111 0.048 0.103 0.064 0.163 0.029 0.056 0.030 0.054 0.042 0.100 0.035 0.082
37.5% 0.034 0.037 0.033 0.060 0.031 0.057 0.035 0.058 0.081 0.180 0.058 0.121 0.057 0.117 0.107 0.229 0.033 0.062 0.032 0.060 0.049 0.111 0.040 0.091
50% 0.031 0.041 0.037 0.065 0.034 0.062 0.038 0.063 0.102 0.207 0.065 0.133 0.066 0.134 0.183 0.312 0.0377 0.068 0.037 0.067 0.053 0.114 0.046 0.099
Avg 0.031 0.036 0.031 0.056 0.030 0.054 0.060 0.144 0.076 0.171 0.055 0.117 0.052 0.110 0.099 0.203 0.032 0.059 0.031 0.057 0.045 0.104 0.038 0.087
Table 10: Full imputation result.
Appendix G Anomaly Detection
Methods MSL SMAP SWaT SMD PSM 1st Pl
Metrics P R F1 P R F1 P R F1 P R F1 P R F1 Count
AnoTransformer 79.61 87.37 83.31 91.85 58.11 71.18 72.51 97.32 83.10 88.91 82.23 85.49 68.35 94.72 79.40 2
Table 11: Anomaly Detection result on 100-length multivariate time series. We calculate Precision, Recall, and F1 score as % for each dataset. ’.’ notation in model name stands for transformer. Bold indicates best result, Underline indicates the second best result.
G.1 SR processing for Anomaly Detection
Spectral Residue
$A(f)=\text{Amplitude}(F(x))$ (18)
$P(f)=\text{Phase}(F(x))$ (19)
$L(f)=\log(A(f))$ (20)
$AL(f)=h_{q}(f)\cdot L(f)$ (21)
$R(f)=L(f)-AL(f)$ (22)
$S(x)=F^{-1}(\exp(R(f)+iP(f)))$ (23)
Appendix H Physics Equations

### v2

@@ 5 Conclusion
In this paper, we introduce TimeDiT, a pioneering approach to creating a versatile and robust foundation model for various time series tasks under practical scenarios. By integrating transformer inductive bias with diffusion model, TimeDiT effectively captures temporal dependencies and addresses real world challenges unique to time series regarding multi-resolution and missing values as well as incorporating external knowledge. Our innovative masking strategies allow for a consistent training framework adaptable to diverse tasks such as forecasting, imputation, and anomaly detection and synthetic data generation. We recognize some limitations of current work: first, we primarily explored common sequence lengths and did not assess TimeDiT’s performance on very long sequences. While we have introduced randomness in prediction length and feature numbers up to a maximum, we aim to develop more scalable solutions for highly variable multivariate time series. Furthermore, our understanding of how different types of domain information contribute to performance improvement is still under investigation. In addition, we acknowledge the importance of sequence-level classification and are actively collecting datasets to extend TimeDiT’s capabilities to classification tasks in future work. Lastly, there is a high demand for deeply developing foundation models for multi-modal time series, allowing TimeDiT to utilize diverse data sources for enhanced performance.
Appendix A TimeDiT Paradigm on Training and Inference
Position of TimeDiT
TimeDiT bridges the gap between ideal foundation models and current practical limitations in time series analysis. While existing approaches often struggle with task-specific architectures or limited domain adaptability, TimeDiT embodies key foundation model characteristics through a balanced design that integrates domain knowledge while maintaining versatile computational capabilities. Its foundation model status is established through three key aspects: (1) a unified mask mechanism that naturally accommodates varying channel sizes and sequence lengths, enabling processing of diverse time series data without task-specific architectures; (2) a versatile framework supporting multiple downstream tasks including forecasting, imputation, anomaly detection, and synthetic data generation; and (3) physics-informed sampling through an energy-based approach that incorporates domain knowledge during inference without model retraining. Although current technical constraints like sequence length limitations and multi-modal integration remain to be fully addressed, this combination of architectural flexibility, task-agnostic design, and domain knowledge integration positions TimeDiT as a practical proto-foundation model that advances the field toward universal time series modeling.
Training details
Similar to the previous DiT work (Peebles and Xie, 2022), TimeDiT is available in four sizes: small (S, 33M parameters), big (B, 130M parameters), large (L, 460M parameters), and extra large (XL, 680M parameters). A comprehensive comparison in Table  10 shows TimeDiT’s expanded task coverage relative to existing general-purpose time series models, including anomaly detection, imputation and data generation. In our training process, we utilized the Adam optimizer with a learning rate of 0.0001 and the loss function is from Equantion 2. Batch sizes of 256 or 512 were employed, depending on model size. The ideal epoch to convergence is over 100 as the complexity of training data, but we choose to use the earlier checkpoint for the case of downstream purpose of anomaly detection and synthetic generation because the two tasks are very dataset-specific and do not necessarily benefit from learning distributions beyond the target dataset. In practice, the maximum channel number ($K_{max}$) was set to 20-40, with a maximum sequence length of 198, unless otherwise specified. All experiments were conducted on NVIDIA A100 GPUs with 40G GPU memory. Importantly, our zero-shot foundation model was trained on Chronos dataset without exposure to any data from the evaluated downstream tasks or datasets. The results of Section 4.1 are derived from a single pre-trained checkpoint, evaluated with or without fine-tuning based on the settings. The only exception is the long-term zero-shot experiments, which require extended sequence inputs while still utilizing the same pre-training dataset. To facilitate reproducibility and further research, we will release the pre-trained checkpoint.
Inference
In the finetuning and inference stage, the choice of mask is tailored to align with the specific requirements of the user. This flexibility allows TimeDiT to apply the most appropriate masking strategy based on the context of the task and application. During inference, while the mask type and parameters are fixed for a given task to ensure consistency, TimeDiT’s generative task architecture allows for flexible transformation of various downstream tasks. This adaptability enables us to address a wide range of time series challenges within a unified framework. Let $n$ represent the number of samples generated for each prediction, which we set to $n=10$ ($n=30$ for forecasting tasks) in our experimental setup at inference time.
Table 9: Comparison of inference times for single-sample generation.
Model Inference Time (mm:ss)
Diffusion-TS 00:06
CSDI 00:02
TimeDiT 00:01
We use the median of these $n$ predictions as the final prediction, providing the added benefit of obtaining a confidence interval for TimeDiT’s predictions. To prevent channel padding from affecting the generated samples, we mask out the invalid channels during sampling at each diffusion timestep so that TimeDiT does not falsely treat the information in the non-valid channels as meaningful information. Padding is applied at the beginning of the temporal dimension to ensure that the most relevant information remains at the end, thereby mitigating the effect of padding. We have included inference time comparisons for single-sample generation, where TimeDiT demonstrates superior computational efficiency, requiring only 1 second for single-sample generation, making it more practical for real-world applications.
Table 10: A comparable analysis of representative general purposes time series models
Model Parameter Size Model Architecture Channel Setting Task Type Pretrain Dataset Data Size
Lag-LLama - Transformer Univariate Forecasting Monash (Godahewa et al., 2021b) 300 Million Time Points
Moriai
S: 14M
B: 91M
L: 311M
Transformer Univariate Forecasting LOTSA (Woo et al., 2024b) 27 Billion Time Points
TimeDiT
S: 33M
B: 130M
L: 460M
XL: 680M
Transformer + Diffusion Multivariate
Forecasting,
Imputation,
Anomaly Detection,
Data Generation
Chronos Ansari et al. (2024) About 5 Billion Time Points
Appendix B Preliminaries of Diffusion Models
In recent years, diffusion models have emerged as a promising approach to generative modeling. A diffusion process is a Markov chain that incrementally adds Gaussian noise to data over a sequence of steps, effectively destroying the data structure in the forward process and reconstructing the data structure during the reverse process.
By iterating this reverse process from $t=T$ down to $t=0$, the model gradually reconstructs the original data from noise. Learning to clean $\mathbf{x}_{T}$ through the reversed diffusion process reduces to building a surrogate approximator to parameterize $\mu_{\theta}(\mathbf{x}_{t},t)$ for all $t$. The reverse process learns to predict the mean and covariance of each intermediate distribution, effectively approximating the original data distribution.
Appendix C Experiments setting
C.1 Datasets
Please refer to Chronos (https://huggingface.co/datasets/autogluon/chronos_datasets) for our pre-trained dataset and the evaluation dataset are listed as follows:
@@ C.1 Datasets
    The ETT (Electricity Transformer Temperature) datasets (Zhou et al., 2021) [^1] ETT: https://github.com/zhouhaoyi/ETDataset include electricity load data at various resolutions (ETTh & ETTm) from two different electricity stations.
@@
    The Weather dataset (Zhou et al., 2021) [^2] Weather:https://www.ncei.noaa.gov/data/local-climatological-data/ comprises 21 meteorological indicators collected in Germany over the span of one year.
@@
    The Electricity (ECL, Electricity Consuming Load) (Zhou et al., 2021) [^3] ECL: https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 dataset provides information on electricity consumption.
@@
    The SMD dataset (Su et al., 2019) includes multivariate time-series data collected from server machines in a data center. It typically contains metrics such as CPU usage, memory usage, and disk activity.
@@
    The PSM dataset (Abdulaal et al., 2021) is used for predictive maintenance and includes sensor data from industrial machines. It often contains readings such as temperature, pressure, and vibration over time.
@@
    The MSL dataset (Hundman et al., 2018) comes from the Mars Science Laboratory mission, specifically the Curiosity rover. It includes telemetry data from the rover’s sensors and systems.
@@
    The SWaT dataset (Mathur and Tippenhauer, 2016) originates from a scaled-down water treatment testbed designed to reflect a real-world water treatment process. It includes sensor and actuator data collected over time.
@@
    The SMAP dataset (Hundman et al., 2018) comes from NASA’s Soil Moisture Active Passive (SMAP) mission, which measures soil moisture and freeze/thaw state. It includes time-series data from multiple sensors aboard the SMAP satellite.
@@
    The Sine dataset (Yoon et al., 2019) is synthetically generated by sinusoidal waves.
@@
    The Air Quality dataset (Yi et al., 2016)  [^4] Air Quality: https://archive.ics.uci.edu/dataset/360/air+qualitycontains hourly averaged readings from five metal oxide chemical sensors integrated into an Air Quality Chemical Multisensor Device. This device was positioned at road level in a highly polluted area of an Italian city. Data were collected from March 2004 to February 2005, making it the longest freely available record of on-field air quality chemical sensor responses.
@@
    The Stock dataset (Yoon et al., 2019) [^5] Stock: https://finance.yahoo.com/quote/GOOG contains daily historical Google stocks data from 2004 to 2019.
@@
    The UCI Appliances Energy prediction dataset (Yoon et al., 2019) [^6] Energy: https://archive.ics.uci.edu/ml/datasetsconsists of multivariate, continuous-valued measurements including numerous temporal features measured at close intervals.
@@
    The Weather_2 dataset (Godahewa et al., 2021a): The Weather_2 dataset comprises hourly climate TSD collected near Monash University, Clayton, Victoria, Australia, from January 2010 to May 2021. It includes series for temperature, dewpoint temperature, wind speed, mean sea level pressure, relative humidity, surface solar radiation, surface thermal radiation, and total cloud cover.
@@
    The PhysioNet dataset (Silva et al., 2012) [^7] The PhysioNet: https://physionet.org/content/challenge-2012/1.0.0/ contains clinical time series data from 12,000 ICU patients, each with 42 vital variables recorded over 48 hours with naturally missing values. Patients are evenly divided into three groups of 4,000 each. For benchmarking purposes, we select 7 out of 42 variables. To address varying scales, we apply standard normalization, resulting in features with zero mean and unit variance.
  15. 15.
    MIMIC-III (Bica et al., 2020) [^8] MIMIC-III: MIMIC-III:https://physionet.org/content/mimiciii/1.4/: MIMIC-III dataset contains 5000 patient ICU records with 19 variables from the lab events table including ‘anion gap, albumin, bands, bicarbonate, bilirubin, creatinine, chloride, glucose, hematocrit, hemoglobin, lactate, platelet, potassium, PTT, INR, PT, sodium, BUN, WBC’. They are irregularly sampled and we process them following the previous works  (Bica et al., 2020; Cao et al., 2023a), which have naturally missing values.
  16. 16.
    NASDAQ: NASDAQ Top 10 Stocks dataset comprises time series data for the ten largest companies by market capitalization listed on the NASDAQ stock exchange. The dataset includes daily and 5-day price data for each stock from 2014-2024, offering two temporal resolutions for comprehensive analysis. We predict the close prices of each company in the multi-resolution forecasting task.
  17. 17.
    Monash dataset archive (Godahewa et al., 2021b): The Monash repository contains 30 datasets, including publicly available time series datasets in various formats and those curated by us. Many datasets have different versions based on frequency and the inclusion of missing values. We use their multivariate time series version for pre-training and evaluation (specified if needed).
@@
PEMS_SF Traffic 4320 852 1 hour
PhysioNet(b) Health Care - 7 Irregular
PhysioNet(b) Health Care - 7 Irregular
PhysioNet(c) Health Care - 7 Irregular
MIMIC-III Health Care - 19 1 day
NASDAQ Finance 2516 20 Multiresolution
C.2 Metrics
@@
$\text{CRPS}(F^{-1},x)\approx\frac{1}{19}\sum_{i=1}^{19}2\Lambda_{i\cdot 0.05}(F^{-1}(i\cdot 0.05),x).$ (19)
@@
CRPS $\text{CRPS Score}=\frac{\sum_{k,l}\text{CRPS}(F^{-1}_{k,l},x_{k,l})}{\sum_{k,l}|x_{k,l}|}$ (20)
where $k$ and $l$ denote the features and time steps of the imputation targets, respectively. The lower the CRPS, the more accurate the model, i.e., the closer the predicted probability is to the observed outcome.
Computations of CRPS_sum
CRPS_sum measures CRPS for the distribution $F$ of the sum of all $K$ features, calculated by:
[…]
@@
Precision
Precision measures the accuracy of positive predictions made by a model. It is defined as the ratio of true positives (TP) to the total number of predicted positives, which includes both true positives and false positives (FP). Mathematically, precision is expressed as:
$\text{Precision}=\frac{TP}{TP+FP}$ (22)
Recall
Recall, also known as sensitivity, measures a model’s ability to correctly identify true positive instances. It is calculated as the ratio of true positives (TP) to the sum of true positives and false negatives (FN). In the context of anomaly detection, failing to detect an anomalous timestamp can have serious consequences, making recall a critical metric. Mathematically, recall is defined as:
$\text{Recall}=\frac{TP}{TP+FN}$ (23)
F1-score
The F1-score is a balanced measure of model performance that combines Recall and Precision. It is calculated as the harmonic mean of these two metrics, giving equal importance to both. This score effectively captures the trade-off between Recall and Precision, penalizing significant disparities between them. By providing a single, comprehensive metric, the F1-score offers a more holistic view of a model’s effectiveness, particularly useful when dealing with imbalanced datasets.
[…]
C.3 Baselines
We conduct a comprehensive comparative analysis, benchmarking TimeDiT against a diverse array of leading models in the field. Our analysis extends to state-of-the-art probabilistic models, encompassing TimeGAN (Yoon et al., 2019), TimeVAE (Desai et al., 2021), Diffusion-TS (Yuan and Qiao, 2024), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021), TransMAF (Rasul et al., 2020), GP-copula (Salinas et al., 2019), and TSDiff (Kollovieh et al., 2023). We also evaluate against cutting-edge deterministic models, including DLinear (Zeng et al., 2023), GPT-2 (Zhou et al., 2023b), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), ETSformer (Woo et al., 2022), FEDformer (Zhou et al., 2022), LightTS (Zhang et al., 2022), Autoformer (Wu et al., 2021), and Anomaly Transformer (Xu et al., 2021), LatentODE and LatentCDE(Rubanova et al., 2019), etc. Furthermore, we include comparisons with recent forecasting foundation models, such as TEMPO (Cao et al., 2023b), Moirai (Woo et al., 2024b), and LagLLama (Rasul et al., 2023). This extensive comparison allows us to thoroughly evaluate TimeDiT’s performance across a wide spectrum of methodologies and architectures in time series modeling.
C.4 Physics Equations in Physics-Informed TimeDiT